OpenStudy (anonymous):

MEDALS!!!! PLEASE HELP!!!!! PLEASE!!!! In the function f(x)=x^2.... How does a decrease of 1 in the independent variable x affect f(x)? A. f(x) will have a change of 1-2x B. f(x) will have a change of x^2-2x+1 C. f(x) will have a change of 2x+1 D. f(x) will have a change of x^2

OpenStudy (anonymous):

@Schen421

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

I am thinking that it may be A

OpenStudy (anonymous):

I'm actually not sure what it's trying to ask, hold on.

OpenStudy (anonymous):

what would happened if you decreased x by 1... i think

OpenStudy (anonymous):

x^2 would be a parabola when graphed, if you were to minus 1 to x^2, then the y-intercept of the graph would shift down by 1 unit

OpenStudy (anonymous):

So it is A?

OpenStudy (anonymous):

Won't be able to give an answer, not too sure about this topic, sorry man!

OpenStudy (anonymous):

Uhhhhhhh.... I have had 5 people try this question.... Hahahaha. This sucks

OpenStudy (anonymous):

do you understand it @Vocaloid ?

Vocaloid (vocaloid):

ok, if x is decreased by one, x becomes (x-1) what happens when you plug (x-1) back into f(x)?

OpenStudy (anonymous):

it will move the parabola 1 unit to the right ?

OpenStudy (anonymous):

I don't know.I really don't understand this

OpenStudy (anonymous):

but i dont the option oO

OpenStudy (anonymous):

see*

OpenStudy (anonymous):

A. f(x) will have a change of 1-2x B. f(x) will have a change of x^2-2x+1 C. f(x) will have a change of 2x+1 D. f(x) will have a change of x^2

OpenStudy (anonymous):

OH IS IT D!?!?!? @Vocaloid

Vocaloid (vocaloid):

no... please do what I asked you to do...

OpenStudy (anonymous):

f(x)-f=x^2 ???

Vocaloid (vocaloid):

no, please substitute (x-1) for x in the equation f(x)

OpenStudy (anonymous):

f(x-1)=x^2

Vocaloid (vocaloid):

no, please check the right side of the equation replace x with (x-1)

OpenStudy (anonymous):

f(x-1)=x-1^2??

Vocaloid (vocaloid):

almost, you're forgetting the parentheses

OpenStudy (anonymous):

f(x-1)=(x-1)^2

Vocaloid (vocaloid):

right! now, try expanding (x-1)^2

OpenStudy (anonymous):

x^2-2x+1

OpenStudy (anonymous):

that would make it B?!

Vocaloid (vocaloid):

great! so, the original equation is f(x) = x^2 the new equation is f(x-1) = x^2 - 2x + 1 what CHANGED? what do f(x) and f(x-1) have in common? what is different?

Vocaloid (vocaloid):

Sorry, no, the answer is not B, please keep trying

Vocaloid (vocaloid):

What makes x^2 different than x^2 - 2x + 1 what part is different about the second equation?

OpenStudy (anonymous):

-2x+1

Vocaloid (vocaloid):

right! so which answer do you think it is?

OpenStudy (anonymous):

C

Vocaloid (vocaloid):

hint: -2x +1 = 1-2x

OpenStudy (anonymous):

I mean A!!

Vocaloid (vocaloid):

great! you got it!!!